A ray of light travelling in air falls on the surface of a transparent material at an angle of 45°to normal. It bends by 15° after refraction. Find the refractive index of the material
1
Expert's answer
2019-06-24T15:46:59-0400

The angle of incidence is 45 degrees.

As the refracted ray bends by 15 degrees, angle of refraction is 45-15=30 degrees.


Using Snell's law:


"n=\frac{\sin{\theta_i}}{\sin\theta_r}=\frac{\sin45^{\circ}}{\sin30^{\circ}}=1.41"

So, the refracting index of the material is n=1.41